how many quarts of a solution with 18% iodine must be added to 40 quarts of water to have a resulting solution with 10% iodine? Can somebody please explain me in details without any Shortcuts:/
Answer by stanbon(75887)

how many quarts of a solution with 18% iodine must be added to 40 quarts of water to have a resulting solution with 10% iodine? Can somebody please explain me in details without any Shortcuts:/
--------------
Equation:
iodine + iodine = iodine
0.18x + 0*40 = 0.10(x+40)
---
Multiply thru by 100 to get:
18x + 0 = 10x + 10*40
8x = 10*40
x = 10*5
x = 50 quarts (amt. of 18% solution needed in the mixture)
========================
Cheers,
Stan H.